a block of mass 10 kg is projected on a rough horizontal surface with a velocity of 10m /sec. calculate how far will it move before coming to rest. coefficient of kinetic friction is 0.1

Grade:12

1 Answers

Arun
6 years ago
Dear Rudra
 

Let 'f' be the coefficient of friction & using ‘g’=10

At final point, the block came to rest,

So at that point force acting in forward direction will be equal to frictional force.

f*mg=m*a

f*g=a

0.10*10=a

a=1

We know that:

v^2 - u^2 = 2as

0-100=2*1*s

s=50

Negative sign appears due to retardation.

 

Regards

Arun (askIITians forum expert)